OpenStudy (xxlyh):

help please?! How many solutions does the equation y + 12 = y + 10 + 2 have?

OpenStudy (xxlyh):

@Lovelarap

OpenStudy (anonymous):

what do you think?

OpenStudy (xxlyh):

I thought it was one but I got it wrong xD

OpenStudy (anonymous):

ok, try to solve it first...

OpenStudy (anonymous):

add the left hand side: y+12=y+12 so.......

OpenStudy (xxlyh):

12+?

OpenStudy (anonymous):

it's no solution: y+12=y+12 solve for y: y-y=12-12 no solution

OpenStudy (xxlyh):

ohhh omg the way I was taught was way different than this but i get it now thank you so much!!

OpenStudy (anonymous):

your welcome! anytime!

DivineSolar (divinesolar):

@xxlyh Close please :).

OpenStudy (unklerhaukus):

y + 12 = y + 10 + 2 y + 12 = y + 12 y = y this is true for any y

OpenStudy (unklerhaukus):

* you are both incorrect so far

DivineSolar (divinesolar):

No. https://gyazo.com/3869ff5b64bb008c788139acd35a4c97

OpenStudy (unklerhaukus):

0 = 0 is true no matter what y is equal to

OpenStudy (unklerhaukus):

for example y could be 3 (3) + 12 = (3) + 10 + 2 15 = 15 or y could be -8 (-8) + 12 = (-8) + 10 + 2 4 = 4 y can be anything

OpenStudy (anonymous):

^ he has point so it would mean all real number?